OpenStudy (anonymous):

The height of a triangle with base b and area A is given by the formula h = . What is the height of a triangle with area 26 cm2 and base 8 cm? A. 1.625 cm B. 3.25 cm C. 3.5 cm D. 6.5 cm

OpenStudy (anonymous):

im thinking b

pooja195 (pooja195):

what formula did you use?

OpenStudy (anonymous):

@freddyfazzbear what postulate did you use?

OpenStudy (anonymous):

idk i forgot

pooja195 (pooja195):

well your answer is incorrect

OpenStudy (anonymous):

i fugured

OpenStudy (anonymous):

Its D.

pooja195 (pooja195):

h=2(26)/8

OpenStudy (anonymous):

Using the formula A=hbb 2 Solving forhb hb=2A b=2·26 8=6.5cm

OpenStudy (anonymous):

medal any1

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

Medal me?

OpenStudy (abhisar):

It's better to give the explanation first and not the answer.

OpenStudy (anonymous):

Look @freddyfazzbear

OpenStudy (anonymous):

just did thx.

OpenStudy (anonymous):

Do you understand it?

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

Okay :)

OpenStudy (anonymous):

ur the best

OpenStudy (anonymous):

Tony wants to calculate his average quiz score. The formula gives Tony’s average score A for a total of T points on n quizzes. Tony had 47 points on 5 quizzes. What is Tony’s average score? points

OpenStudy (anonymous):

ok so do 47/5?

pooja195 (pooja195):

9.4 and thats what u get when u do divide it

pooja195 (pooja195):

:) good job

OpenStudy (anonymous):

thx.

OpenStudy (anonymous):

The formula d = rt gives the distance traveled in time t at rate r. A delivery truck drives for 1.5 hours and travels 76 miles. What was the average speed of the truck? A. 38 miles per hour B. 50.7 miles per hour C. 74.5 miles per hour D. 114 miles per hour

OpenStudy (anonymous):

do 1.5 times 76?

pooja195 (pooja195):

yep

OpenStudy (anonymous):

i got 114

pooja195 (pooja195):

wait yoou divide

OpenStudy (anonymous):

i timesed

pooja195 (pooja195):

76/1.5

OpenStudy (anonymous):

The formula A = l • w gives the area of a rectangle. A rectangle has an area of 37.2 cm2 and a length of 12 cm. What is the width of the rectangle? cm